OpenStudy (anonymous):

Confused, 4/3 pi(2r)^3 . It says to simplify , I don't know how the book got it but the answer was 32/3 pi(r^3)

OpenStudy (anonymous):

please write it so we can see it more accurately. what is raised to the 3rd power? use equation below the post.

OpenStudy (anonymous):

ok sorry gove me a sec

OpenStudy (anonymous):

\[4/3\pi(2r)^3 \] is supposed to be simplified

OpenStudy (anonymous):

\[(2r)^{3} = 2^{3}* r ^{3} = 8*r ^{3}\] \[8 * \frac{4}{3} = \frac{32}{3}\] so we insert the pi back in and get \[\frac{32}{3} * \Pi * r ^{3}\]

OpenStudy (anonymous):

Oh jeeze, that seemed more simpler than I thought. Thanks. I love this community